Find the equation of a line that passes through the point (1,2) and has a slope of 1/3. Show how you found the equation.
My book doesn't show how to solve with a fraction involved....
This is my attempt at the equation. y=1/3x+2 is this correct?

Answer by algebrahouse.com(1659)

This is one way to do it:
(1,2), m = 1/3

y - y1 = m(x - x1) is point-slope form

y - 2 = 1/3(x - 1) {substituted point and slope into point-slope form}
y - 2 = 1/3x - 1/3 {used distributive property}
y = 1/3x + 5/3 {added 2 to both sides}
www.algebrahouse.com